# RPNT: Robust Pre-trained Neural Transformer - A Pathway for Generalized Motor Decoding

### v1

@@ 4.1 Datasets and experimental setup
Experimental setup We have conducted experiments of RPNT function and validation first focused on self-supervised pretraining of the RPNT model and then evaluation and comparison of RPNT downstream decoding generalization with baselines through supervised finetuning (SFT) on test recordings (see Figure 1E for model adaptation). For benchmark (Perich et al., 2018), we followed the baseline setup according to (Azabou et al., 2023; Ryoo et al., 2025), where we used the train session data in subjects C, J, and M for pretraining and tested with the rest of the sessions for downstream evaluations. Specifically, our decoding evaluation included the following three cases: 1) New center-out sessions from Monkey C (denoted as C-CO) for cross-session evaluation; 2) New center-out sessions from Monkey T (denoted as T-CO) for cross-subject evaluation; 3) New random target sessions from Monkey T (denoted as T-RT) for cross-subject and cross-tasks evaluation. Notably, more than $80\%$ of sessions in the benchmark are CO tasks. Thus, the T-RT evaluation is expected to be the most challenging task, since it requires the model to adjust to both subject and task distribution shift. We followed the released code and customized it for building the train/valid/test datasets according to the Python package torch brain  (Azabou et al., 2023). The details for the processing are also seen in the Appendix E. For the neuropixel dataset, we used 16 (S1-S16) sites of recording neural activity (with 0.8/0.1/0.1 for train/val/test split) to pretrain the RPNT model, then used the remaining site (S17) (with 0.2/0.3/0.5 for train/val/test split) for FS cross-site evaluation. Further, as the CO task was highly structured, we followed the convention by only evaluating the decoding performance at the reaching period on the successful trials (Ryoo et al., 2025). We used $R^{2}$ to quantify the decoding performance between the predicted and the ground-truth velocity behaviors. The baseline comparison models are described in the appendix I.
4.2 Cross-session, cross-subject, and cross-task evaluation results on public benchmark
@@
From scratch Wiener filter 0.8712 $\pm$ 0.0137 0.6597 $\pm$ 0.0392 0.5942 $\pm$ 0.0564
MLP 0.9210 $\pm$ 0.0010 0.7976 $\pm$ 0.0220 0.7007 $\pm$ 0.0774
@@
indicate results are copied from the POSSM (Ryoo et al., 2025) paper. We only select the FT: Full model finetuning results that are directly related to our experimental setup. For more comprehensive table results can check their paper directly. Acknowledgment is included in the appendix I.
@@
Table 1: Velocity decoding performance ($R^{2}$) comparison across three generalization scenarios (C-CO, T-CO, and T-RT) on the public benchmark. RPNT (FS-SFT) used few-shot finetuning, whereas RPNT (Full-SFT) used all available downstream sessions. The mean and standard deviation were calculated across all sessions. Bold numbers indicated that the SOTA performance in each scenario. The references for baseline models are listed in the appendix I.
@@
We compare the velocity decoding performance ($R^{2}$) across three generalization scenarios on the benchmark in Table 1, where RPNT was evaluated with two types of baselines and training: 1) models trained from scratch on single downstream sessions and 2) pretrained models finetuned with different amounts of downstream sessions.
@@
Training from scratch. In this regime (top segment of Table 1), compared models are trained for each downstream session only. RPNT achieved the best performance across all three scenarios. Specifically, RPNT obtained $R^{2}$ scores (mean $\pm$ std) of $0.9647\pm 0.0026$, $0.9103\pm 0.0182$, and $0.8356\pm 0.0914$ in C-CO, T-CO, and T-RT evaluations, respectively. These results show enhancement with respect to existing decoding models, with particularly notable achievement in the challenging T-RT scenario ($\approx 7\%$ increase from second second-best baseline POSSM-GRU).
Pretrained models. We next investigated the benefits of pretraining and a downstream supervised finetuning (SFT) approach. As a self-supervised pretraining model, RPNT finetuning can operate in two regimes: a) Finetuning on the full training dataset of the downstream task (Full-SFT) and b) More restrictive few-shot finetuning, where finetuning is performed on a single downstream session (FS-SFT). RPNT was tested in these two regimes and compared with baselines in the FS-SFT regimes. Results (bottom segment of Table 1) indicate that both RPNT(Full-SFT) and RPNT(FS-SFT) outperformed previous baselines in the FS-SFT regime. Notably, the more challenging regime of testing RPNT, RPNT(FS-SFT), achieved average $R^{2}$ scores of $0.9801$, $0.9431$, and $0.8515$ for three tasks, in contrast to existing state-of-the-art POYO and POSSM baselines in the FS-SFT regime, which had access to behavior labels for all sessions during the pretraining phase. In the comparable label usage of Full-SFT, RPNT $R^{2}$ was further enhanced compared to the results on the FS-SFT regime and reached $0.9894\pm 0.0037$ (C-CO), $0.9626\pm 0.0059$ (T-CO), and $0.8778\pm 0.1005$ (T-RT). Due to these results, and the potential of the FS-SFT regime to require more efficient finetuning in latter experiments of section 4.3 and section 4.4, we fixed RPNT in the regime of FS-SFT. As we report in the results below, RPNT surpasses the performance of existing models, indicating the efficiency and robustness of the RPNT in decoding task generalization. With respect to RPNT variants, we note that our results indicate that RPNT that is pretrained outperforms RPNT that is trained from scratch, e.g., Pretrained RPNT (Full-SFT): 0.8778 vs. RPNT from scratch 0.8356 in the T-RT task, indicating the benefits of the self-supervised pretraining strategy for robust neural representation learning.
4.3 Cross-site evaluation result on the Neuropixel dataset
@@
Wiener filter 0.3462 $\pm$ 0.0710
MLP 0.4074 $\pm$ 0.0592
RNN (LFADS) 0.5015 $\pm$ 0.1085
Transformer (NDT) 0.5272 $\pm$ 0.0720
Transformer (PoYo) 0.5944 $\pm$ 0.0901
@@
Figure 3: Decoding performance comparison across sites (B-CS) on the neuropixel dataset. The mean and standard deviation were calculated across all successful trials.
@@
Figure 4: Sweep analysis for training splits.
@@
We further investigated the generalization of RPNT for variations that include cross-site scenarios on the neuropixel dataset. This dataset is distinct from the public benchmark in recording modality (neuropixels vs. microelectrodes) and is particularly suitable for cross-site variation study since these recordings did not share neurons across sites (see appendix figure 2 for recording sites topology). Figure 4 summarizes the decoding performance of RPNT compared with existing baselines. RPNT (in FS-SFT regime) outperforms existing baselines at the downstream single-site (S17) velocity decoding task. Furthermore, pretraining RPNT improved performance (scratch: 0.6358 vs. pretrained: 0.6612). In addition, we swept train splits from $1\%$ to $50\%$ ($20\%$ by default), while keeping the test data to be the last $50\%$. We observe that even with $10\%$ of training split (half of the prior), pretrained RPNT is able to achieve reasonable performance (see Figure 4), which shows the potential for few-shot model tuning applications. A comprehensive study of sweeping rain split for all models listed in Figure 4 can be found in the appendix table 11.
4.4 Ablation Studies
Method T-RT B-CS
Sinusoidal PE 0.8260 0.6242
RoPE 0.8226 0.6484
Learnable PE 0.8305 0.6273
MRoPE 0.8515 0.6612
Table 2: Ablation study of position embeddings
Method T-RT B-CS
Standard attention 0.8024 0.5024
Context-based attention 0.8515 0.6612
Table 3: Ablation study of attention

### v2

@@ Results
We conducted experiments to evaluate RPNT, first focusing on self-supervised pretraining and then on downstream decoding generalization through supervised finetuning (SFT) with baseline comparisons on test recordings (see Figure 1E for model adaptation). In LTRCH [40], we followed the setup in POSSM [43]. We used designated training sessions from subjects C, J, and M for pretraining, and the remaining sessions for downstream evaluation. Our decoding evaluation included three cases: 1) New center out sessions from Monkey C (denoted as C-CO) for cross session evaluation; 2) New center-out sessions from Monkey T (denoted as T-CO) for cross subject evaluation; 3) New random target sessions from Monkey T (denoted as T-RT) for cross subject and cross tasks evaluation. Since more than $80\%$ of sessions in the LTRCH benchmark are center-out tasks, T-RT is the most challenging evaluation, since it requires adaptation to both subject and task distribution shifts. We followed the released code and customized it to build the train/validation/test datasets according to the Python package torch brain [1]. Details of the processing are provided in the Appendix 1. For NPCS, we used 16 sites (S1-S16) for pretraining (0.8/0.1/0.1 for train/validation/test split), and the remaining site (S17) was used for Few-Shot (FS) SFT of Cross Site (CS) evaluation (0.2/0.3/0.5 train/validation/test split). As the CO task is highly structured, we evaluated decoding performance during the reaching period on successful trials following prior work [43]. For RT tasks, we evaluated the performance using the whole period for all trials. We used $R^{2}$ to quantify decoding performance between predicted and ground truth velocity. Baseline comparison models are listed in the Appendix 5.
Cross-session, cross-subject, and cross-task evaluation results on LTRCH
@@
From scratch Wiener filter [54] 0.8712 $\pm$ 0.0137 0.6597 $\pm$ 0.0392 0.5942 $\pm$ 0.0564
MLP †, [22] 0.9210 $\pm$ 0.0010 0.7976 $\pm$ 0.0220 0.7007 $\pm$ 0.0774
@@
These results were obtained and reported in recent POSSM work [43] and cited in this table for comparison. The pretrained section includes the Full Model Finetuning (FT) results [43], which are directly related to the RPNT experimental setup.
@@
Table 1: Velocity decoding performance ($R^{2}$) comparison across three generalization scenarios (C-CO, T-CO, and T-RT) on the LTRCH benchmark. RPNT (FS-SFT) uses few shot finetuning, whereas RPNT (Full-SFT) uses all available downstream sessions. Mean and standard deviation are computed across sessions. Bold numbers indicate the best performance in each scenario. References for baseline models are provided in Appendix 5 and comments and .
@@
We compared velocity decoding performance ($R^{2}$) across three generalization scenarios on the LTRCH benchmark (Table 1). RPNT was evaluated under two training settings: (1) models trained from scratch on single downstream sessions and (2) pretrained models finetuned with different amounts of downstream data.
@@
Training from scratch. We first trained models from scratch using a single downstream session. We compared previously published models with RPNT following the training and evaluation approach used in recent work (POSSM) [43] (Table 1, top). RPNT achieved the best performance across all three generalization challenges (cross-session, cross-subject, and cross-task). RPNT obtained $R^{2}$ scores (mean $\pm$ std) of $0.9647\pm 0.0026$, $0.9103\pm 0.0182$, and $0.8356\pm 0.0914$ in C-CO, T-CO, and T-RT evaluations, respectively. These results demonstrated improved performance over existing decoding models, with a notable gain in the challenging T-RT scenario (approximately $7\%$ over the second-best baseline POSSM-GRU).
Pretrained models. We next investigated the benefits of model pretraining paired with downstream supervised finetuning (SFT) and compared performance in this regime across models. As a self-supervised pretraining model, we evaluated RPNT in two regimes: (a) finetuning on the full downstream training dataset (Full-SFT), and (b) few-shot finetuning on a single downstream session (FS-SFT). While we evaluated RPNT in both regimes, we compared its performance with that of other previously published models evaluated in the FS-SFT setting using existing evaluations (provided in POSSM [43]; Table 1, bottom). Our results showed that both RPNT (Full-SFT) and RPNT (FS-SFT) outperformed existing baselines, obtaining average $R^{2}$ scores of $0.9801$, $0.9431$, and $0.8515$. This was despite the fact that both RPNT regimes did not have any access to behavioral labels at any stage (both pretraining and testing), while existing baselines (e.g., POYO [1] and POSSM [43]) in the FS-SFT setting had access to behavior labels from all sessions during pretraining. This shows that RPNT performs well in the computationally efficient and more constrained training settings imposed by the FS-SFT regime. We therefore focused solely on this setting in subsequent experiments (Section Cross-site evaluation result on the Neuropixel dataset and Section Ablation studies). RPNT Performance was higher in the Full-SFT regime, compared to FS-SFT, producing $R^{2}$ scores (mean $\pm$ std) 0.9894 $\pm 0.0037$ (C-CO), $0.9626\pm 0.0059$ (T-CO), and $0.8778\pm 0.1005$ (T-RT) across tasks, showing that RPNT was also able to leverage access to multiple sessions to improve generalization. Overall, we found that RPNT consistently surpassed existing models, demonstrating improved efficiency and robustness in decoding generalization.
Cross-site evaluation result on the Neuropixel dataset
@@
Wiener filter [54] 0.3462 $\pm$ 0.0710
MLP [22] 0.4074 $\pm$ 0.0592
RNN (LFADS) [37] 0.5015 $\pm$ 0.1085
Transformer (NDT) [64] 0.5272 $\pm$ 0.0720
Transformer (PoYo) [1] 0.5944 $\pm$ 0.0901
@@
Table 2: Decoding performance comparison across sites (B-CS) on the neuropixel dataset. We train the baseline models from scratch on the NPCS dataset. The mean and standard deviation were calculated across all successful trials.
@@
Figure 3: Sweep analysis for training splits for pretrained RPNT model.
@@
We further investigated the generalization of RPNT under Cross Site variations using the NPCS dataset. This dataset differs from the LTRCH benchmark in recording modality (Neuropixels vs. microelectrodes) and is particularly suitable for cross site analysis, as recordings do not share neurons across sites (see Figure 2 for recording site topology). We compared decoding performance on RPNT with previously published model architectures with available implementation code (Table 2). RPNT, in the FS-SFT regime, outperformed all existing benchmark model architectures on the downstream single site (S17) velocity decoding task. Furthermore, pretraining consistently improved performance (scratch: 0.6358 vs. pretrained: 0.6612).
We also swept the amount of data used for training (training split) from $1\%$ to $50\%$ ($20\%$ is set as default) while fixing the amount of data used for testing (test split) to $50\%$. We observed that even with $10\%$ training data (half of the default $20\%$ FS-SFT train split), the pretrained RPNT model still maintained reasonable performance ( Figure 3, slightly above 0.5 $R^{2}$), demonstrating its potential uses for applications with limited downstream data. RPNT outperformed other benchmark models tested across the majority of training splits assessed, with the exception of the extreme low-data regime of $1\%$ (Table 3).
Method 1% 5% 10% 20% 30% 40% 50%
Wiener filter 0.0009 0.0344 0.0892 0.3462 0.3567 0.3597 0.3634
MLP 0.1181 0.3070 0.3704 0.4074 0.4215 0.4279 0.4314
RNN -0.0049 -0.0215 0.1083 0.5015 0.5534 0.6229 0.6295
NDT -0.0038 0.0836 0.2401 0.5272 0.5752 0.6082 0.6175
PoYo -0.0136 0.2285 0.5378 0.5944 0.6578 0.6938 0.7285
RPNT (pretrained) 0.0701 0.3735 0.5418 0.6612 0.6927 0.7170 0.7329
[…]
Transfer datasets performance of our pretrained RPNT model
Method C-CO T-CO T-RT
Transfer from NPCS to LRTCH 0.8603 $\pm$ 0.0009 0.7585 $\pm$ 0.0344 0.6395 $\pm$ 0.0892
No transfer, only on LRTCH 0.9894 $\pm$ 0.0037 0.9626 $\pm$ 0.0059 0.9626 $\pm$ 0.0059
Table 4: Evaluation results of the NPCS pretrained RNPT model on the LRTCH dataset
Method B-CS
Transfer from LRTCH to NPCS 0.6537 $\pm$ 0.0413
No transfer, only at NPCS 0.6612 $\pm$ 0.0328
Table 5: Evaluation results of the LRTCH pretrained RNPT model on the NPCS dataset
Both datasets consist of nonhuman primates performing similar motor tasks with recordings from similar brain regions, which gave us the opportunity to perform two experiments to evaluate transfer generalization of pretrained RPNT across datasets. First, we applied the NPCS pretrained RPNT model to the LTRCH dataset. We observed that RPNT had limited generalization performance (Table 5). In contrast, when we reversed the transfer direction by applying RPNT pretrained on the LTRCH dataset to the NPCS dataset, we observed that RPNT maintained strong generalization performance (transfer: 0.6537 vs. original: 0.6612) (Table 5). These asymmetric results suggested that pretraining on the LTRCH dataset enabled more robust neural representations that could transfer effectively to the NPCS Neuropixel dataset. One possible explanation was the difference in data scale, where the LTRCH dataset contained more than 100 sessions, in contrast with 16 sessions on the NPCS dataset. These results further suggested the scalability benefits of leveraging larger datasets during the pretraining phase of RPNT.
Ablation studies
The RPNT model architecture contains multiple novel components. We next evaluated how each component contributed to RPNT’s decoding performance using ablation studies. We conducted ablations of the positional encoding, attention mechanism, and masking strategy evaluated on the T-RT and B-CS tasks in the FS-SFT regime (Tables 8, Tables 8, and Tables 8). We performed additional sweeping analyses (e.g., transformer layers and attention heads), summarized in the Appendix ( 6).
Positional encoding. We compared RPNT’s positional encoding approach, MRoPE, with three alternative positional encoding approaches: sinusoidal PE, standard RoPE, and multi-dimensional learnable PE (Table 8). MRoPE achieved the highest decoding performance ($R^{2}$ score). An improvement of approximately $3\%$ over standard RoPE supported the extension of RoPE to multiple dimensions, enabling generalization to session configurations while preserving relative position information.
Context-based Attention mechanism. We compared RPNT’s context-based attention design with standard self-attention mechanisms (Table 8). This comparison demonstrated the importance of the context-based attention design, as we observe a substantial performance drop of approximately $5\%$ when shifting to standard self-attention. This gap highlights the effectiveness of the adaptive kernel mechanism in handling the nonstationary nature of neural recordings. We further include a sweep study over different kernel sizes, in which we observed that context-based attention also consistently performed better than the standard attention. (see Appendix Table 7).
Random uniform masking strategy. We compared RPNT’s random uniform masking strategy to multiple fixed masking ratio strategies, including symmetric and asymmetric across neuron and temporal dimensions. All fixed masking ratio strategies consistently underperformed the random uniform masking strategy (Table 8). This result suggested that exposing the model to a range of masking ratios leads to more robust neural representations and improved downstream performance.
Together, these ablation studies confirmed that each proposed component contributed to the strong performance of RPNT in generalized motor decoding.
Method Cross-Task (T-RT) Cross-Site (B-CS)
Sinusoidal PE 0.8260 $\pm$ 0.0894 0.6242 $\pm$ 0.0267
RoPE 0.8226 $\pm$ 0.0941 0.6484 $\pm$ 0.0074
Learnable PE 0.8305 $\pm$ 0.0917 0.6273 $\pm$ 0.0398
MRoPE 0.8515 $\pm$ 0.1071 0.6612 $\pm$ 0.0328
Table 6: Ablation study of positional encoding methods.
Method Cross-Task (T-RT) Cross-Site (B-CS)
Standard attention 0.8024 $\pm$ 0.0896 0.5024 $\pm$ 0.0388
Context-based attention 0.8515 $\pm$ 0.1071 0.6612 $\pm$ 0.0328
Table 7: Ablation study of attention mechanisms.